OpenStudy (anonymous):

(9-2i)(3+i) simplify

OpenStudy (anonymous):

Hint: Use Distributive law \[i^2 = -1\]

Parth (parthkohli):

Just like your algebra...just like your algebra

OpenStudy (shubhamsrg):

how about de-moiver's theorem ? just to make it interesting? ;)

OpenStudy (anonymous):

\[(9-2i)(3+i)=27+9i-6i-2i^2=27+3i+2=29+3i\]